Questions 21 through 25 refer to the following output: ∑x = 210        ∑x2 = 2870     ∑xy =...

  1. Questions 21 through 25 refer to the following output:

    ∑x = 210        ∑x2 = 2870     ∑xy = -10745

    ∑y = -759   ∑y2 = 62523

    Regression Equation:  = 5.873684 – 4.173684x

    Source

    SS

    df

    MS

    F

    Regression

    11548.06053

    1

    11584.06053

    9.42011

    Error

    22134.88947

    18

    1229.716

    Total

    33718.95

    19

    21. How many observations were used in the above regression?

1 points   

QUESTION 22

  1. What percentage of the observed variability in Y is explained by X?

1 points   

QUESTION 23

  1. What is the correlation between X and Y?

1 points   

QUESTION 24

  1. Give the value of the t statistic for testing whether the slope between X and Y is zero.

1 points   

QUESTION 25

  1. What's the predicted value of Y when X is 9?
Answer #1

21) observations were used in the above regression =19+1 =20

22) percentage of the observed variability in Y is explained by X =r2*100

=( 11548.06053/ 33718.95 )*100 =34.25%

23) orrelation between X and Y = -0.5852

24) value of the t statistic = -sqrt(F)= -3.069

25)

predicted value = 5.873684 – 4.173684*9=-31.6895